# Selection of Filters for Photonic Crystal Spectrometer Using Domain-Aware Evolutionary Algorithms

### v1

@@ 3.2 Design of the Objective Function
This section aims to propose a criterion to estimate the quality of the candidate solution. Let us denote the set of concentrations of different gases as $\mathscr{C}$ and the set of all other possible configurations of TGMD as $\mathscr{R}$. Example of configurations in $\mathscr{R}$ are parameters of the exact gas concentration retrieval algorithm. The noise in the atmosphere is modeled as an object $\xi$ taken from the set $\Omega$. The particular instance $\xi$ is taken from the space $\Omega$ by chance every time the simulation of the gas measurement is made. In our approximation of gas receiving by TGMD, we consider the environment where TGMD is located to be homogeneous, meaning that noise does not depend on the time of the made measurement. Therefore the probability measure $P\left(\left\{\xi\right\}\right)$ is chosen in advance and does not depend on time for all $\xi\in\Omega$. When all the parameters of the gas measurement device are fixed, it is modeled as a function
@@
In the context of our application, we consider only methane as a gas to detect, so the set $\mathscr{C}$ contains only different concentrations of methane. Therefore, in our case $\mathscr{C}\subset\mathbb{R}$. Since object $\xi\in\Omega$ is taken by chance, the function $\mathrm{S}$ defines a stochastic process. The distribution of each random variable does not have to coincide with distribution $P$ due to the complicated process of gas restoring described in Section 2. For the convenience of the reader, we gather all notation used here in Table 1.
@@
$N$ $640$, the length of the sequence of filters;
$M$ $16$, the dimensionality of candidate solutions to OFS;
@@
$\mathrm{D}(\bm{x},\xi)$ stochastic process defined in Eq. (3) that assess the quality of solution $\bm{x}$ to OFS problem;
@@
[…]
As described in Section 2, we aim to locate the best filters $\bm{x}\in\mathscr{L}$ for the single gas concentration $c^{*}\in\mathscr{C}$ and single configuration of the system $\bm{r}^{*}\in\mathscr{R}$. Therefore, we assume $c^{*}$ and $\bm{r}^{*}$ fixed and so simplify the notation to define a function $F$ that grades the performance of chosen $\bm{x}\in\mathscr{L}$. This function is defined on the space $\mathscr{L}$ and returns a real value from $\mathbb{R}$, which denotes the error made in gas retrieval.
@@
Informally, filters $\bm{x}$ should be graded better than $\bm{y}$ by the function $F$ if $\bm{x}$ leads to the same or better precision in the restored concentration of gas as $\bm{y}$, but with a smaller “uncertainty”.
This requirement 1 is motivated by a wish to obtain a solution that allows for precise detection of the gas with strong confidence. Now we will give a more formal interpretation of our intuitive requirement.
@@
3.2.1 Properties of the Defined Objective Function
Let us show the properties of this objective function. First of all, it conforms to our intuitive criteria 1 for the assessment of filter selections. Indeed, for the fixed solutions $\bm{x},\bm{y}\in\mathscr{L}$ the requirement can be rewritten as follows:
$\begin{cases}\operatorname{E}^{2}\left[\mathrm{D}(\bm{x},\xi)\right]\leq\operatorname{E}^{2}\left[\mathrm{D}(\bm{y},\xi)\right]\\ \operatorname{Var}\left[\mathrm{D}(\bm{x},\xi)\right]<\operatorname{Var}\left[\mathrm{D}(\bm{y},\xi)\right].\end{cases}$ (5)
Let us demonstrate how function $F$ influences the additional difference in means needed to compensate for the increase in variances. Consider some fixed selections of filters $\bm{x},\bm{y}\in\mathscr{L}$ that satisfy the Eq. (6). Due to the symmetry between $\bm{x}$ and $\bm{y}$, we assume without loss of generality that:
$\mathrm{F}(\bm{x})\leq\mathrm{F}(\bm{y})$ (7)
Case 1).
Case 2).
$\dfrac{\operatorname{E}^{2}(\mathrm{Y})}{\operatorname{Var}(\mathrm{Y})}\cdot\left(\operatorname{E}^{2}(\mathrm{Y})-\operatorname{E}^{2}(\mathrm{X})\right)\geq\\ \dfrac{\operatorname{E}^{2}(\mathrm{Y})}{\operatorname{Var}(\mathrm{Y})}\cdot\left(\operatorname{Var}(\mathrm{X})-\operatorname{Var}(\mathrm{Y})\right).$ (9)
Refer to caption Refer to caption
Case 3).
Analysis.
We compute the sample mean of the set $\mathbb{V}$, which gives an approximation of the value in LHS of Eq. (10) and Eq. (11) expected while optimizing the function $F$. The $\ln$ of this value equals $-6.6$, and it is depicted by the gray plane in Figure 6. The part of the surface that is lower than the gray plane in Figure 6 depicts the set of ratios of expected values and means that lead to grading $\mathrm{X}$ better than $\mathrm{Y}$ in average over the set $\mathbb{V}$. Conversely, the region above the plane represents ratios leading to worse average grading of $\mathrm{X}$ compared to $\mathrm{Y}$.
When $\operatorname{Var}(\mathrm{X})$ grows relative to $\operatorname{Var}(\mathrm{Y})$, the value ${\operatorname{Var}(\mathrm{X})}/{\operatorname{Var}(\mathrm{Y})}$ becomes bigger. As demonstrated in Figure 6, in this situation it becomes less likely to asses $\mathrm{X}$ better than $\mathrm{Y}$. The intersection between the surface and the gray plane in Figure 6 occurs when the ratio of variances, $\operatorname{Var}(\mathrm{X})/\operatorname{Var}(\mathrm{Y})$, is less than $1.0024$. This intersection depicts the maximum tolerable ratio in variances where $\mathrm{X}$ can still be graded better than $\mathrm{Y}$ by the function $F$, on average over the set $\mathbb{V}$. When such grading is possible, we see that it becomes less likely when values of $\operatorname{E}(\mathrm{X})$ and $\operatorname{E}(\mathrm{Y})$ are closer in absolute value. Moreover, it is clear that for the fixed ratio of expected value the RHS of Eq. (10) is monotonically increasing, so grading of $\mathrm{X}$ better than $\mathrm{Y}$ under the assumption in Eq. (6) is possible only when variances are not too different but means compensate this difference by far. In summary, we demonstrated how the means must adjust to offset the rise in variances. Although the results of this analysis are not directly applied in the subsequent sections of the paper, they remain crucial as they provide guarantees regarding the candidate solutions we favor.
@@ 3.2.2 Computation of the Objective Function on a Given Solution
It is not possible to compute moments of random variable $\mathrm{D}^{2}(\bm{x})$ precisely due to the unconstrained nature of the simulation $\mathrm{S}$. In this regard, we estimate the value of $\mathrm{F}(x)$ using the sample mean. Specifically, we sample $K$ times the value of random variable $\mathrm{D}^{2}(\bm{x},\xi)$ and compute an approximation $\operatorname{\widehat{E}}_{K}\left[\mathrm{D}^{2}(\bm{x},\xi)\right]$ of $\operatorname{E}\left[\mathrm{D}^{2}(\bm{x},\xi)\right]$ by finding the arithmetical average of the sampled values. Then, the approximation of $\mathrm{F}(\bm{x})$ is
@@
This is a realistic assumption because the distributions of random variables $\mathrm{D}(\bm{x},\xi)$ are close to normal for several $\bm{x}$ that we considered, for example, see Figure 17. Therefore, we can reasonably assume that the distribution of random variable $\mathrm{D}^{2}(\bm{x},\xi)$ is not heavy tail for every $\bm{x}\in\mathscr{L}$.
@@
[…]
@@
When, for some $\bm{x},\bm{y}\in\mathscr{L}$, the difference between $\widehat{\mathrm{F}}_{K_{1}}(\bm{x})$ and $\widehat{\mathrm{F}}_{K_{2}}(\bm{y})$ becomes small, we can use Welch’s unequal variances t-test [32] to reject or accept the null hypothesis that the estimated means are different. This test imposes the constraint that the sample means of random variables are normally distributed, which is satisfied in our case due to the chosen large value of $K$ in Eq. (13). This test was chosen because 1) we do not assume that the variances of the compared random variables are equal, and 2) the test allows us to compare populations with different numbers of samples, meaning $K_{1}\neq K_{2}$, which we find convenient for implementation. We apply this test in Algorithm 4 described in Section 5.3.
@@ 4 Optimization of OFS problem
A representative selection of Meta-heuristics and Machine Learning approaches is considered to solve the OFS formulated in Section 3.1 in the Problem Formulation 1 with the objective function $\mathrm{F}$ defined in Section 3.2. The choice of these numerical methods is justified for the reasons discussed in Section 1. We discuss the particular selected algorithms and then compare the results of their application to OFS in Section 4.5. Sections 4.1, 4.2, 4.3 correspond to the classes of algorithms defined in [23].
Obviously, we do not argue that we considered all the possible methods since it is infeasible to account for every numerical approach. We acknowledge the possibility that there might be other optimization algorithms, not considered in this study, that could potentially identify even better solutions to OFS within the given computational limits. However, the chosen algorithms demonstrate the effectiveness of applying optimization techniques to our OFS problem. Notably, several algorithms outperform the baseline approach described in Section 2. This success establishes the feasibility of efficient OFS optimization within the defined time constraints and highlights the potential for obtaining high-quality solutions. Future work can investigate both unaddressed optimization algorithms and more advanced versions of the ones already considered. This focus on further improvement becomes even more relevant as we plan to enhance the realism of our simulator, leading to more computationally expensive simulations.
@@ 4.1 Evolutionary Algorithms
Randomized search heuristics (RSH) are off-the-shelf algorithms to solve complex optimization problems [33, 34]. Such algorithms define the distribution over candidate solutions, and sample populations of candidates from the distribution and usually adjust the distribution based on the qualities of the sampled candidates. Evolutionary Algorithms [34, 35] are nature-inspired RSH that use metaphors of evolution. These algorithms produce descendant candidate solutions by a sequence of reproductions, selections, recombinations, and mutations. The qualities of those solutions are evaluated by the objective function.
The pseudo-boolean search space, meaning set $\left\{0,1\right\}^{n}$ for fixed constant $n\in\mathbb{N}$, is usually considered in the context of discrete optimization. $(1+1)$ EA is an Evolutionary Algorithm, which was rigorously analyzed on several pseudo-boolean toy problems, for example, see [36, 37]. The algorithm maintains a candidate solution $\mathbf{x}_{\text{best}}\in\{0,1\}^{n}$ with the smallest (in the case of minimization) observed value of the objective function $F$. The optimization process works in multiple iterations. At every iteration of the discussed $(1+1)$ EA, a number $k$ is sampled from Binomial Distribution $\mathcal{B}(n,r/n)$. Here $n$ is the size of the vector, which represents the candidate solution, and $r\in(0,n]$ is a real constant, which represents the mutation rate. If $k=0$, then we change it to 1 (so-called “shift” strategy), since it is more efficient for practical purposes [38]. Then $k$ components in the solution $\mathbf{x}_{\text{best}}$ are chosen uniformly at random. A new candidate solution $\mathbf{x}_{\text{new}}$ is created by flipping the bits in each of the $k$ chosen components. The new candidate solution $\mathbf{x}_{\text{new}}$ is taken as new observed solution if and only if $F(\mathbf{x}_{\text{best}})\geq F(\mathbf{x}_{\text{new}})$. Such iterations repeat until the computation budget is exhausted, or other termination conditions are satisfied.
@@ 4.1.1 $(\mu+\lambda)$ EA
In the present work, we consider one adaptation of $(1+1)$ EA to the integer domains proposed in [39]. Specifically, the mutation of every component is modified for integer domains. Assume that every component is bounded and can contain integers from $1$ to $c$. Then we change the integer in every selected component to an integer chosen uniformly at random from $1$ to $c$. Notably, mutation, when the newly chosen integer is bound to be different from the existing integer in the component, was called uniform mutation by the authors of [39]. However, we omitted this detail in our implementation. We did not check the equality between the newly chosen integer and the existing value during mutation. In the context of OFS problem, this simplification is unlikely to impact the performance significantly. The probability of mutating to the existing value remains low due to the large value of the constant $c$ since it is equivalent to $L$, the number of filters in the library. The scheme to select the number of mutated components is maintained the same as in $(1+1)$ EA.
$(\mu+\lambda)$ EA explores the implementation of this algorithm with a bigger population size for solving OFS. In this algorithm:
@@
  • Selection and Mutation: during each iteration $\lambda$ offspring are created. Each offspring is generated in two steps: 1) a single parent solution is selected uniformly at random from the current population of $\mu$ candidates; 2) The mutation operation described earlier is applied to this parent solution.
  • Elitist Survival Selection: The algorithm then combines the parent population ($\mu$ candidates) with the newly generated $\lambda$ offspring ($\mu+\lambda$ total solutions). From this combined pool the algorithm selects the $\mu$ best individuals (those with the smallest objective function values) to form the population for the next iteration. This selection process ensures that high-performing solutions are carried forward.
In our implementation, we used the constants $\mu=10,\lambda=20$.
@@ 4.1.2 $(\mu/2+\lambda)$ EA
Apart from mutations, evolutionary algorithms use crossovers to generate new candidate solutions [40]. In this work, we consider uniform crossover which produces new candidate solutions by combining two given solutions. Specifically, given solutions $\mathbf{x},\mathbf{y}$ uniform crossover generates solution $\mathbf{z}$ such that every component $z_{i}$ is set to either $x_{i}$ or $y_{i}$ with equal probability. Paired Crossover Evolutionary Algorithm [41] is an algorithm that uses only uniform crossover and selections. It proved itself as one of the algorithms capable of addressing practical SCOPs [42, 43]. However, all solutions generated by this algorithm consist of only the integers that occurred in the initial population. It means that, when applied to OFS problem, this algorithm is limited in the exploration of all the possible combinations of filters from the library. Therefore, we use uniform crossover in combination with mutations and embed it in the structure of $(\mu+\lambda)$ EA discussed in 4.1.1.
$(\mu/2+\lambda)$ EA selects two candidate solutions uniformly at random from the population at every iteration. The uniform crossover is applied to the selected solutions to produce the new candidate solution. Then the same mutation is applied to this created solution. The type of mutation and the rest of the implementation details remain the same as in $(\mu+\lambda)$ EA. The notation $\mu/2$ in the name of the algorithm emphasizes that two parents take part in the uniform crossover.

### v2

@@ 3.2 Design of the Objective Function
This section proposes a criterion to estimate the quality of the candidate solution. Let us denote the set of concentrations of different gases as $\mathscr{C}$ and the set of all other possible configurations of TGMD as $\mathscr{R}$. Example of configurations in $\mathscr{R}$ are parameters of the exact gas concentration retrieval algorithm. Atmospheric noise is modeled by a random variable $\xi$ drawn from a noise space $\Omega$. A new instance of $\xi$ is sampled from $\Omega$ each time a gas measurement simulation is performed, reflecting the stochastic nature of atmospheric variability. In our approximation of gas receiving by TGMD, we consider the environment where TGMD is located to be homogeneous, meaning that noise does not depend on the time of the measurement. Therefore, the probability measure $P\left(\left\{\xi\right\}\right)$ is chosen in advance and does not depend on time for all $\xi\in\Omega$. When all the parameters of the gas measurement device are fixed, it is modeled as a function
@@
In the context of our application, we consider only methane as a gas to detect, so the set $\mathscr{C}$ contains only different concentrations of methane. Therefore, in our case $\mathscr{C}\subset\mathbb{R}$. Since object $\xi\in\Omega$ is taken by chance, the function $\mathrm{S}$ defines a stochastic process. For the convenience of the reader, we gather all notation used here in Table 1.
@@
NN $640$, the length of the sequence of filters;
$M$ $16$, the dimensionality of candidate solutions to OFS, which is the number of selected filter types;
@@
$\mathrm{D}(\bm{x},\xi)$ stochastic process defined in Eq. (2) that assesses the quality of solution $\bm{x}$ to OFS problem;
@@
$\textsc{u.a.r.}\!\left(\mathbb{X}\right)$ select an element uniformly at random from set $\mathbb{X}$;
$\textsc{u.a.r.}\!\left(a,b\right)$ select a real value uniformly at random from the segment $a,b]\cap\mathbb{R}$.
As described in Section 2, we aim to locate the best filters $\bm{x}\in\mathscr{L}$ for the single gas concentration $c^{*}\in\mathscr{C}$ and single configuration of the system $\bm{r}^{*}\in\mathscr{R}$. Therefore, we assume $c^{*}$ and $\bm{r}^{*}$ fixed and so simplify the notation to define a function $F$ that qualifies the performance of the chosen $\bm{x}\in\mathscr{L}$. This function is defined on the space $\mathscr{L}$ and returns a real value from $\mathbb{R}$, which denotes the error made in gas retrieval.
@@
Informally, filter set $\bm{x}$ should be qualified better than $\bm{y}$ by the function $F$ if $\bm{x}$ leads to the same or better precision in the retrieved concentration of gas as $\bm{y}$, but with a smaller “uncertainty”.
This requirement is motivated by a wish to obtain a solution that allows for precise detection of the gas with strong confidence. Now we will give a more formal interpretation of our intuitive requirement.
@@
The proposed OFS objective Eq. (3) both meets our intuitive requirements and admits an analytic characterization of the mean-variance trade‑off. We propose an analysis of its properties in the A, and here we keep only the conclusions. The function compromises the variance and mean of the retrieved gas concentration. In the conflicting case of these objectives, we propose the necessary and sufficient conditions for qualifying a solution as better than others. Together, these results guarantee how much increase in mean is required to compensate any rise in variance, providing justification for the behavior of $F$ in selecting certain filter combinations.
@@ 3.2.1 Computation of the Objective Function on a Given Solution
It is not possible to compute moments of random variable $\mathrm{D}^{2}(\bm{x})$ precisely due to absence of the analytical form of the simulation $\mathrm{S}$. In this regard, we estimate the value of $\mathrm{F}(x)$ using the sample mean. Specifically, we sample $K$ times the value of random variable $\mathrm{D}^{2}(\bm{x},\xi)$ and compute an approximation $\operatorname{\widehat{E}}_{K}\left[\mathrm{D}^{2}(\bm{x},\xi)\right]$ of $\operatorname{E}\left[\mathrm{D}^{2}(\bm{x},\xi)\right]$ by finding the arithmetical average of the sampled values. Then, the approximation of $\mathrm{F}(\bm{x})$ is
@@
This is a realistic assumption because the distributions of random variables $\mathrm{D}(\bm{x},\xi)$ are close to normal for several $\bm{x}$ that we considered, for example, it becomes clear later in Figure 14. Therefore, we can reasonably assume that the distribution of random variable $\mathrm{D}^{2}(\bm{x},\xi)$ is not heavy-tailed for every $\bm{x}\in\mathscr{L}$.
@@
$\widehat{\mathrm{F}}_{K}(\bm{x})\xrightarrow{d}\mathcal{N}\!\left(\operatorname{E}\!\left[\mathrm{D}^{2}(\bm{x},\xi)\right],\dfrac{\operatorname{Var}\!\left[\mathrm{D}^{2}(\bm{x},\xi)\right]}{K}\right).$ (5)
@@
When, for some $\bm{x},\bm{y}\in\mathscr{L}$, the difference between $\widehat{\mathrm{F}}_{K_{1}}(\bm{x})$ and $\widehat{\mathrm{F}}_{K_{2}}(\bm{y})$ becomes small, we use Welch’s unequal variances t-test [28] to reject or accept the null hypothesis that the estimated means are different. This test imposes the constraint that the sample means of random variables are normally distributed, which is satisfied in our case due to the chosen large value of $K$ in Eq. (5). This test was chosen because 1) we do not assume that the variances of the compared random variables are equal, and 2) the test allows us to compare populations with different numbers of samples, meaning $K_{1}\neq K_{2}$, which we find convenient for implementation. We apply this test in Algorithm 4 described in Section 5.4.
@@ 4 Optimization of OFS problem
A representative selection of metaheuristics and Machine Learning approaches is considered to solve the OFS formulated in Section 3.1 in the Problem Formulation 1 with the objective function $\mathrm{F}$ defined in Section 3.2. The choice of these numerical methods is justified for the reasons discussed in Section 1. We discuss the particular selected algorithms and then compare the results of their application to OFS in Section 4.5. Sections 4.14.24.3 correspond to the classes of algorithms defined in [19], which provides a comprehensive review of simulation optimization algorithms and their applications, ensuring that our selection of methods is grounded in established best practices within the field.
We acknowledge the possibility that there might be other optimization algorithms, not considered in this study, that could potentially identify even better solutions to OFS within the given computational limits. However, the chosen algorithms demonstrate the effectiveness of applying optimization techniques to our OFS problem. Notably, several algorithms outperform the baseline approach described in Section 2. This success establishes the feasibility of efficient OFS optimization within the defined time constraints and highlights the potential for obtaining high-quality solutions. Future work can investigate both unaddressed optimization algorithms and more advanced versions of the ones already considered. This focus on further improvement becomes even more relevant as we plan to enhance the realism of our simulator, leading to more computationally expensive simulations.
@@ 4.1 Evolutionary Algorithms
Randomized search heuristics (RSH) are off-the-shelf algorithms to solve complex optimization problems [29, 30]. Such algorithms define the distribution over candidate solutions, sample populations of candidates from the distribution and usually adjust the distribution based on the qualities of the sampled candidates. Evolutionary Algorithms [30, 31] are nature-inspired RSH that use metaphors of evolution. These algorithms produce descendant candidate solutions by a sequence of reproductions, selections, recombinations, and mutations. The quality of those solutions is evaluated by the objective function.
The pseudo-boolean search space $\left\{0,1\right\}^{n}$ for fixed constant $n\in\mathbb{N}$ is usually considered in the context of discrete optimization. $(1+1)$ EA is an Evolutionary Algorithm, which was rigorously analyzed on several pseudo-boolean toy problems, for example, see [32, 33]. The algorithm maintains a candidate solution $\mathbf{x}_{\text{best}}\in\{0,1\}^{n}$ with the smallest (in the case of minimization) observed value of the objective function $F$. The optimization process works in multiple iterations. At every iteration of the discussed $(1+1)$ EA, a number $k$ is sampled from Binomial Distribution $\mathcal{B}(n,r/n)$. Here $n$ is the size of the vector, which represents the candidate solution, and $r\in(0,n]$ is a real constant, which represents the mutation rate. If $k=0$, then we change it to 1 (so-called “shift” strategy), since it is more efficient for practical purposes [34]. Then $k$ components in the solution $\mathbf{x}_{\text{best}}$ are chosen uniformly at random. A new candidate solution $\mathbf{x}_{\text{new}}$ is created by flipping the bits in each of the $k$ chosen components. The new candidate solution $\mathbf{x}_{\text{new}}$ is taken as new observed solution if and only if $F(\mathbf{x}_{\text{best}})\geq F(\mathbf{x}_{\text{new}})$. Such iterations repeat until the computation budget is exhausted, or other termination conditions are satisfied.
@@ 4.1.1 $(\mu+\lambda)$ EA
In this paper we study a variant of the $(1+1)$ EA for integer variables as proposed in [35]. Each variable can take any integer value from 1 to $c$. During mutation, a selected variable is replaced by an integer chosen uniformly at random from 1 to $c$. In [35] this operator is called uniform mutation and it includes a check that the new value must differ from the old one. In our implementation we skip that check. This means we allow the new value to be the same as the old one. Since $c$ is large and equals the number of filters $L$, the probability of drawing the same value twice is very small. Apart from this simplification, we follow the standard $(1+1)$ EA rule for choosing how many variables to mutate.
$(\mu+\lambda)$ EA is an implementation of this algorithm with a bigger population size for solving OFS. In this algorithm:
@@
  • Selection and Mutation: during each iteration $\lambda$ offspring are created. Each offspring is generated in two steps: 1) a single parent solution is selected uniformly at random from the current population of $\mu$ candidates; 2) The uniform mutation operation described in this section is applied to this parent solution.
  • Elitist Survival Selection: The algorithm then combines the parent population ($\mu$ candidates) with the newly generated $\lambda$ offspring ($\mu+\lambda$ total solutions). From this combined pool the algorithm selects the $\mu$ best individuals (those with the smallest objective function values) to form the population for the next iteration.
To justify our design choices in the $(\mu+\lambda)$ EA, we note the following points. We choose a population of size $\mu$ and generate $\lambda$ offspring in each generation to balance exploration and computational cost. A larger population improves diversity and reduces the risk of premature convergence. Selecting each parent uniformly at random maintains an unbiased search, meaning that it avoids over‐emphasising any individual. We apply elitist survival selection ensures that the best solutions are never lost and that the algorithm makes steady progress toward better filters for the OFS problem.
In our implementation, we used $\mu=10,\lambda=20$.
@@ 4.1.2 $(\mu/2+\lambda)$ EA
Apart from mutations, evolutionary algorithms use crossover to generate new candidate solutions [36]. In this work, we consider uniform crossover which produces new candidate solutions by combining two given solutions. Specifically, given solutions $\mathbf{x},\mathbf{y}$ uniform crossover generates solution $\mathbf{z}$ such that every component $z_{i}$ is set to either $x_{i}$ or $y_{i}$ with equal probability. The Paired Crossover Evolutionary Algorithm [37] relies solely on uniform crossover and selection. It has strong theoretical guarantees on simple test functions and it works well on various SCOPs [38, 39].
However, this algorithm only uses the values that appeared in the initial population. In the OFS problem, we need to explore new combinations of filters from a large library. To overcome this limitation, we add our mutation operator to the procedure. We embed uniform crossover and mutation in the $(\mu+\lambda)$ EA framework from Section 4.1.1.
In our $(\mu/2+\lambda)$ EA, we pick two parents at random from the current population of size $\mu$. We apply uniform crossover to these parents to produce one child. Then we propose to apply the mutation described in Section 4.1.1. We repeat this process $\lambda$ times each iteration to generate $\lambda$ offspring. We combine the $\mu$ parents with the $\lambda$ offspring and select the best $\mu$ solutions to form the next population. The notation $\mu/2$ highlights that two parents are involved in each crossover step.